.Let A = {1, 2, 3, 4, 5} and B = {0, 3, 6}. Find a) A ∪ B. b) A ∩ B. c) A − B. d) B − A


1
Expert's answer
2022-03-27T17:29:52-0400

"A\cup B\,\,is\,\,the\,\,union, i.e. all\,\,elements\,\,of\,\,A\,\,and\,\,all\,\,elements\,\,of\,\,B:\\A\cup B=\left\{ 0,1,2,3,4,5,6 \right\} \\A\cap B\,\,is\,\,the\,\,inter\sec tion, i.e. all\,\,elements\,\,which\,\,are\,\,both\,\,in\,\,A\,\,and\,\,in\,\,B:\\A\cap B=\left\{ 3 \right\} \\A-B\,\,consists\,\,of\,\,all\,\,elements\,\,which\,\,are\,\,in\,\,A\,\,except\,\,for\,\,those\,\,which\,\,are\,\,in\,\,B:\\A-B=\left\{ 1,2,4,5 \right\} \\B-A\,\,consists\,\,of\,\,all\,\,elements\,\,which\,\,are\,\,in\,\,B\,\,except\,\,for\,\,those\,\,which\,\,are\,\,in\,\,A:\\B-A=\left\{ 0,6 \right\}"